# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ 4.2. Polynomial conditions for the existence of spaces $L_{\lambda p^{n-1},n}$
In terms of the $Q_{i}$’s, we have from Remark 4.3 that
$\frac{P}{P_{n}}=(-1)^{n-1}\prod_{\underline{\epsilon}\in{\mathbb{F}}_{p}^{n-1}}(Q_{n}+\epsilon_{n-1}Q_{n-1}+\epsilon_{n-2}Q_{n-2}+\dots+\epsilon_{1}Q_{1}),$
which in particular implies that the coefficients of $\left(\frac{P}{P_{n}}\right)^{p-1}$ are polynomial expressions in the coefficients of the $Q_{i}$’s. In what follows, we aim to describe these polynomial expressions more precisely. We begin by stating and proving the following proposition, which is independent of previous results obtained in this paper. This result is interesting also because it leads to a direct proof (i.e. not relying on the case $n=2$) of Theorem 4.7.
Proposition 4.19.
Let $I_{n}:={\mathbb{F}}_{p}^{n}$. For every element $\underline{\epsilon}\in I_{n}$, let $s(\underline{\epsilon})=\sum_{i=1}^{n}\epsilon_{i}$. If
$A:=\prod_{\underline{\epsilon}\in I_{n-1}}\left(X_{n}+\epsilon_{n-1}X_{n-1}+\epsilon_{n-2}X_{n-2}+\dots+\epsilon_{1}X_{1}\right)^{p-1}$
and
$B:=\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n}\\ s(\underline{\epsilon})\neq 0\end{subarray}}\left(\epsilon_{n}X_{n}+\epsilon_{n-1}X_{n-1}+\epsilon_{n-2}X_{n-2}+\dots+\epsilon_{1}X_{1}\right),$
then
$A+B=\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})\neq-1\end{subarray}}\left(X_{n}+\epsilon_{n-1}X_{n-1}+\epsilon_{n-2}X_{n-2}+\dots+\epsilon_{1}X_{1}\right)^{p}.$
In particular, we have that $A\equiv-B\mod k[X]^{p}$.
Proof.
Let us write $B=B_{1}B_{2}$ with
$B_{1}:=\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n}\\ s(\underline{\epsilon})\neq 0,\epsilon_{n}\neq 0\end{subarray}}\left(\epsilon_{n}X_{n}+\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)\;\mbox{and}\;B_{2}=\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})\neq 0\end{subarray}}\left(\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right).$
Let us show that the polynomial $B_{1}$ contains the common factors between $A$ and $B$: we can rewrite this polynomial as
$B_{1}$ $=\prod_{\epsilon_{n}\in{\mathbb{F}}_{p}^{\times}}\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})\neq-\epsilon_{n}\end{subarray}}\left(\epsilon_{n}X_{n}+\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)$
$=\prod_{\epsilon_{n}\in{\mathbb{F}}_{p}^{\times}}\epsilon_{n}^{(p-1)p^{n-2}}\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})\neq-\epsilon_{n}\end{subarray}}\left(X_{n}+\frac{\epsilon_{n-1}}{\epsilon_{n}}X_{n-1}+\dots+\frac{\epsilon_{1}}{\epsilon_{n}}X_{1}\right)$
$=\prod_{\epsilon_{n}\in{\mathbb{F}}_{p}^{\times}}\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})\neq-1\end{subarray}}\left(X_{n}+\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)$
$=\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})\neq-1\end{subarray}}\left(X_{n}+\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)^{p-1},$
and this shows that $A+B=B_{1}C$ with
$C:=\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})=-1\end{subarray}}\left(X_{n}+\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)^{p-1}+B_{2}.$
In order to conclude, we need to show that
$C=\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})\neq-1\end{subarray}}\left(X_{n}+\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right),$
which is equivalent to consider $C$ as a monic univariate polynomial of degree $(p-1)p^{n-2}$ in the variable $X_{n}$ and show that its set of roots is $\{\delta_{n-1}X_{n-1}+\dots+\delta_{1}X_{1}:\underline{\delta}\in I_{n-1},s(\underline{\delta})\neq 1\}.$ To verify this, for every $\underline{\delta}\in I_{n-1}$ with $s(\underline{\delta})\neq 1$ we set $c:=s(\underline{\delta})-1$ and we substitute $X_{n}$ with $\delta_{n-1}X_{n-1}+\dots+\delta_{1}X_{1}$ in the expression of $C-B_{2}$. This gives
$\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})=-1\end{subarray}}\left(\delta_{n-1}X_{n-1}+\dots+\delta_{1}X_{1}+\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)^{p-1}=\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})=c\end{subarray}}\left(\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)^{p-1}=$
$=\prod_{\epsilon_{n}\in{\mathbb{F}}_{p}^{\times}}\left(\epsilon_{n}^{p^{n-2}}\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})=c\end{subarray}}\big{(}\frac{\epsilon_{n-1}}{\epsilon_{n}}X_{n-1}+\dots+\frac{\epsilon_{1}}{\epsilon_{n}}X_{1}\big{)}\right)=\prod_{\epsilon_{n}\in{\mathbb{F}}_{p}^{\times}}\left(\epsilon_{n}\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})=\frac{c}{\epsilon_{n}}\end{subarray}}\left(\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)\right)$
$=-\prod_{\begin{subarray}{c}\underline{\epsilon}\in I_{n-1}\\ s(\underline{\epsilon})\neq 0\end{subarray}}\left(\epsilon_{n-1}X_{n-1}+\dots+\epsilon_{1}X_{1}\right)=-B_{2}.$
As a result, $\delta_{n-1}X_{n-1}+\dots+\delta_{1}X_{1}$ is a zero of $C\in k[X_{1},\dots,X_{n-1}][X_{n}]$for every $\underline{\delta}\in I_{n-1}$ satisfying $s(\underline{\delta})\neq 1$. ∎
Remark 4.20.
As a corollary of Proposition 4.19 we get a new proof of Theorem 4.7, which is of quite a different nature than the one given in Section 4.1. More precisely, employing the same notation as in the Theorem, the non-trivial part is to show that, if the form $\omega_{n}$ is logarithmic, then the forms $\omega_{i}$ are logarithmic for every $i=1,\dots,n$. We then need to prove that, if $\left(\frac{P}{P_{n}}\right)^{p-1}$ satisfies the system of equations (4.18) then $\left(\frac{P}{P_{i}}\right)^{p-1}$ also satisfies (4.18) for every $i$. In fact, we can apply Remark 4.3 to show that $\left(\frac{P}{P_{i}}\right)^{p-1}$ is, up to a sign, equal to $A(Q_{i},Q_{1},\dots,\widehat{Q_{i}},\dots,Q_{n})$. From this, it follows that $\left(\frac{P}{P_{i}}\right)^{p-1}$ can be obtained from $\left(\frac{P}{P_{n}}\right)^{p-1}$ by applying a permutation of the $Q_{i}$’s.
On the other hand, Proposition 4.19 shows that the coefficients of $A$ appearing in equations (4.18) are the same as those of $-B$, which is symmetric in the $Q_{i}$’s, hence we have that $\left(\frac{P}{P_{n}}\right)^{p-1}$ satisfies (4.18) if, and only if $\left(\frac{P}{P_{i}}\right)^{p-1}$ satisfies (4.18).
4.2.1. Many new examples of spaces $L_{\lambda 2^{n-1},n}$
As an application of Proposition 4.19, we construct for all values of $n\geq 2$ new large classes of examples of spaces $L_{\lambda 2^{n-1},n}$ in characteristic $2$. In this case, we have that $p-1=1$, and the equations (4.18) are equivalent to the condition
$A(Q_{1},\dots,Q_{n})-X\in k[X]^{2}.$
We can describe the coefficients of $A(Q_{1},\dots,Q_{n})$ appearing in (4.18) in more detail.
Proposition 4.21.
Let $p=2$ and let ${\mathfrak{A}}_{n}$ be the alternating group on $n$ letters. Then, we have
$A(X_{1},\dots,X_{n})\equiv\sum_{\sigma\in{\mathfrak{A}}_{n}}X_{\sigma(1)}X_{\sigma(2)}X_{\sigma(3)}^{2}\cdots X_{\sigma(n)}^{2^{n-2}}\mod\;k[X]^{2}.$
Proof.
The polynomial $A$ is homogeneous of degree $2^{n-1}$, and is additive in the variable $X_{n}$. By Proposition 4.19, its reduction modulo $k[X]^{2}$ is invariant with respect to any permutation of the $X_{i}$’s, and hence it is additive in $X_{i}$ for every $i$. As a result, the monomials appearing in $A$ mod $k[X]^{2}$ are all of the form $\prod_{i=1}^{n}\alpha_{i}X_{i}^{a_{i}}$ with $a_{i}\in\{2^{j},0\leq j\leq 2^{n-2}\}$ satisfying $\sum_{i=1}^{n}a_{i}=2^{n-1}$. Since we are studying only the terms that are not squares, at least one of the $a_{i}$’s needs to be equal to 1. We claim that the conditions above uniquely determine the exponents $a_{i}$. More precisely, let us assume for simplicity that we have $a_{1}\leq\dots\leq a_{n}$. We claim that then we have $a_{1}=1$, $a_{2}=1$, $a_{3}=2$, $a_{4}=2^{2}$, …, $a_{n}=2^{n-2}$.
Proof of the claim.
Since the $a_{i}$’s are powers of $2$, in order to show the claim it is enough to show that $a_{i}\leq 2^{i-2}$ for all $i=2,\dots,n$. We can do this by finite induction on $i$: we certainly need to have $a_{2}=1$ since the sum of all the terms is even. We then assume that, for every $j\leq i$, we have $a_{j}\leq 2^{j-2}$. It follows that
$2^{n-1}-a_{i+1}-\dots-a_{n}=a_{1}+\dots+a_{i}\leq 2^{i-1}.$
Applying the $2$-adic valuation $v_{2}$, we have that
$v_{2}(2^{n-1}-a_{i+1}-\dots-a_{n})=v_{2}(a_{1}+\dots+a_{i})\leq i-1.$
Applying the non-archimedean triangular inequality gives
$v_{2}(2^{n-1}-a_{i+1}-\dots-a_{n})=v_{2}(a_{i+1}+\dots+a_{n})\geq v_{2}(a_{i+1}),$
which allows us to conclude that $v_{2}(a_{i+1})\leq j-1$, that is $a_{i+1}\leq 2^{i-1}$. ∎
As a result of the claim, the monomials appearing in $A$ mod $k[X]^{2}$ are precisely the required ones. We now have to show that the coefficients are all equal to 1. By symmetry, it is enough to show that the coefficient of $\big{(}\prod_{i=1}^{n-2}X_{i}^{2^{n-i-1}}\big{)}X_{n-1}X_{n}$ in the polynomial $A$ is equal to 1. To produce such a term, we pick $X_{1}$ from all the $2^{n-2}$ factors where it appears (those with $\epsilon_{1}=1$), we pick $X_{2}$ from the $2^{n-3}$ factors where $X_{2}$ appears (those with $\epsilon_{1}=0$, $\epsilon_{2}=1$), and so on, until we remain with the expression $X_{n}(X_{n-1}+X_{n})$ from which we need to pick the monomial $X_{n-1}X_{n}$. Since this is the only process that produces the monomial $\big{(}\prod_{i=1}^{n-2}X_{i}^{2^{n-i-1}}\big{)}X_{n-1}X_{n}$, its coefficient in $A$ is equal to 1. ∎
Corollary 4.22.
Let $p=2$ and let $\underline{Q}\in k[X]^{n}$ be a $n$-uple giving rise to a basis of a space $\Omega$ as in Definition 4.4. Then $\Omega$ is a space $L_{\lambda 2^{n-1},n}$ if, and only if
$\det(\underline{Q^{\prime}},\underline{Q},\underline{Q^{2}},\dots,\underline{Q}^{2^{n-2}})=1.$
Proof.
As a direct consequence of Proposition 4.21, we have that the system of equations (4.18) is equivalent to $\left(\sum_{\sigma\in{\mathfrak{A}}_{n}}Q_{\sigma(1)}Q_{\sigma(2)}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}\right)-X\in k[X]^{2},$ which is equivalent to
$\left(\sum_{\sigma\in{\mathfrak{A}}_{n}}Q_{\sigma(1)}Q_{\sigma(2)}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}\right)^{\prime}=1.$
By denoting $\tau\in{\mathfrak{S}}_{n}$ the transposition exchanging 1 and 2, we can express the left hand side of this equation in the desired determinantal form. In fact, we have that
$\big{(}\sum_{\sigma\in{\mathfrak{A}}_{n}}Q_{\sigma(1)}Q_{\sigma(2)}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}\big{)}^{\prime}=\sum_{\sigma\in{\mathfrak{A}}_{n}}Q_{\sigma(1)}^{\prime}Q_{\sigma(2)}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}+\sum_{\sigma\in{\mathfrak{A}}_{n}}Q_{\sigma(1)}Q_{\sigma(2)}^{\prime}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}$
$=\sum_{\sigma\in{\mathfrak{A}}_{n}}Q_{\sigma(1)}^{\prime}Q_{\sigma(2)}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}+\sum_{\sigma\in{\mathfrak{A}}_{n}}Q_{\sigma\tau(1)}^{\prime}Q_{\sigma\tau(2)}Q_{\sigma\tau(3)}^{2}\cdots Q_{\sigma\tau(n)}^{2^{n-2}}$
$=\sum_{\sigma\in{\mathfrak{A}}_{n}}Q_{\sigma(1)}^{\prime}Q_{\sigma(2)}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}+\sum_{\sigma\in{\mathfrak{S}}_{n}-{\mathfrak{A}}_{n}}Q_{\sigma(1)}^{\prime}Q_{\sigma(2)}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}$
$=\sum_{\sigma\in{\mathfrak{S}}_{n}}Q_{\sigma(1)}^{\prime}Q_{\sigma(2)}Q_{\sigma(3)}^{2}\cdots Q_{\sigma(n)}^{2^{n-2}}=\det(\underline{Q^{\prime}},\underline{Q},\underline{Q^{2}},\dots,\underline{Q}^{2^{n-2}}).$
If $n=2$, Corollary 4.22 leads to a complete classification of spaces $L_{2\lambda,2}$ in characteristic 2. In fact, there are unique polynomials $U_{1},U_{2},V_{1},V_{2}\in k[X]$ such that $Q_{1}=U_{1}^{2}+XV_{1}^{2}$ and $Q_{2}=U_{2}^{2}+XV_{2}^{2}$. With these notations, we have $Q_{1}^{\prime}=V_{1}^{2}$ and $Q_{2}^{\prime}=V_{2}^{2}$, and hence
The condition imposed by Corollary 4.22 then is equivalent to
$V_{1}U_{2}+U_{1}V_{2}=1.$
If $\lambda$ is odd, then the polynomials $V_{1},V_{2}$ have degree $\frac{\lambda-1}{2}$. By Bézout’s theorem, if $V_{1}$ and $V_{2}$ are coprime, then there exists a unique pair $(U_{1},U_{2})$ with $\deg(U_{i})<\frac{\lambda-1}{2}$ satisfying the above condition. If we pick $(V_{1},V_{2})$ such that $V_{1}+V_{2}$ has also degree $\frac{\lambda-1}{2}$, and we let $(U_{1},U_{2})$ be the pair given by Bézout’s theorem, then by Corollary 4.22 the polynomials $U_{1}^{2}+XV_{1}^{2}$ and $U_{2}^{2}+XV_{2}^{2}$ give rise to a basis of a space $L_{2\lambda,2}$. If we relax the condition $\deg(U_{i})<\frac{\lambda-1}{2}$, then we have other pairs that satisfy Bezout’s theorem: these can be obtained from the minimal one $(U_{1},U_{2})$ as $(U_{1}+aV_{1},U_{2}+aV_{2})$ for $a\in k$.
If $\lambda$ is even, then the argument above works by choosing coprime $U_{1},U_{2}\in k[X]$ of degree $\frac{\lambda}{2}$ with $U_{1}+U_{2}$ of degree $\frac{\lambda}{2}$ and applying Bézout’s theorem to find $V_{1},V_{2}$. This shows that spaces $L_{2\lambda,2}$ exist for every $\lambda$ and in large abundance. While this is already known by work of Pagot [7, Théorème 2.2.4], the approach relying on Corollary 4.22 is of a different nature.
For $n>2$, this approach is in principle not enough to give a complete classification, but we can extract sufficient conditions for the existence of spaces $L_{2^{n-1}\lambda,n}$ that lead to the discovery of new large classes of examples for every $n$.
Proposition 4.23.
Let $p=2$ and $n\geq 3$. Let $\underline{Q}\in k[X]^{n}$ be a $n$-uple giving rise to a basis of a space $\Omega$ as in Definition 4.4. For every $i=1,\dots,n$ let $U_{i},V_{i}\in k[X]$ be such that $Q_{i}=U_{i}^{2}+XV_{i}^{2}$ and that they satisfy the system of equations
(4.24) $\begin{cases}\det(\underline{U},\underline{V},\underline{U}^{2},\dots,\underline{U}^{2^{n-2}})=1&\\ \det\left(\underline{U},\underline{V},((1+\epsilon_{3})\underline{U}+\epsilon_{3}\underline{V})^{2},\dots,((1+\epsilon_{n})\underline{U}+\epsilon_{n}\underline{V})^{2^{n-2}}\right)=0,&\end{cases}$
@@ 4.2.1. Many new examples of spaces $L_{\lambda 2^{n-1},n}$
We need to show that, if $\underline{U}$ and $\underline{V}$ satisfy (4.24), then $\underline{Q}$ satisfy the condition of Corollary 4.22. This latter is equivalent to $\det\left(\underline{V}^{2},\underline{U}^{2},\underline{Q^{2}},\dots,\underline{Q}^{2^{n-2}}\right)=1$, which is in turn equivalent to
@@
If the system (4.24) is satisfied, then the non-constant coefficients of the polynomial above vanish and the constant term is equal to 1. Hence we can apply Corollary 4.22 to get that $\Omega$ is a space $L_{\lambda 2^{n-1},n}$. ∎
@@
  • We then check the equations in the second line of (4.24): for each $(\epsilon_{3},\dots,\epsilon_{n})\in{\mathbb{F}}_{2}^{n-2}-\{0\}$ we let $k\in\{3,\dots,n\}$ be the smallest number such that $\epsilon_{k}\neq 0$. Since we need every $\epsilon_{i}$ with $i<k$ to be equal to 0, the vectors $\left(\underline{U},\underline{V},((1+\epsilon_{3})\underline{U}+\epsilon_{3}\underline{V})^{2},\dots,((1+\epsilon_{k})\underline{U}+\epsilon_{k}\underline{V})^{2^{k-2}}\right)$ forming the first $k$ columns inside the determinant can be rewritten as $\left(\underline{U},\underline{V},\underline{U}^{2},\dots,\underline{U}^{2^{k-1}},\underline{V}^{2^{k-2}}\right)$. By condition $(i)$, we have that all these column vectors belong to the $k-1$ dimensional space generated by $\underline{V},\dots,\underline{V}^{2^{k-2}}$ and hence
    $\det\left(\underline{U},\underline{V},\underline{U}^{2},\dots,\underline{U}^{2^{k-1}},\underline{V}^{2^{k-2}},\dots,((1+\epsilon_{n})\underline{U}+\epsilon_{n}\underline{V})^{2^{n-2}}\right)=0.$
@@
Let $\underline{W}:=\leavevmode\nobreak\ (W_{1},\dots,W_{n})\subset k[X]^{n}$ be a $n$-uple of polynomials such that all the nonzero elements of $\langle W_{1},\dots,W_{n}\rangle_{{\mathbb{F}}_{2}}$ are pairwise coprime and of the same degree, denoted by $d$ (in particular, we have that $\Delta_{n}(\underline{W})\neq 0$). If we set $V_{i}:=\Delta_{n-1}(\underline{\hat{W_{i}}})$ for every $i=1,\dots,n$, then there exists a rational function $\alpha\in\frac{1}{\Delta_{n}(\underline{W})}k[X]$ satisfying the following properties:
@@
    The $n$-uple $\underline{Q}=(Q_{1},\dots,Q_{n})\in k[X]^{n}$ gives rise to a basis of a space $L_{2^{n-1}\lambda,n}$.
@@
We begin by observing that the hypothesis that $V_{i}=\Delta_{n-1}(\underline{\hat{W_{i}}})$ for every $i$, combined with Theorem A.3 implies that
@@
We claim that this system is equivalent to the condition that $P_{{\mathfrak{W}}_{i}}(W_{i})$ divides $\gamma+V_{i}$ for every $i\in\{1,\dots,n\}$. In fact, let $i\neq j$ and set ${\mathfrak{W}}_{ij}:={\mathfrak{W}}_{i}\cap{\mathfrak{W}}_{j}$. If $W\in\langle W_{1},\dots,W_{n}\rangle-\{0\}$ is a common divisor of $P_{{\mathfrak{W}}_{i}}(W_{i})$ and $P_{{\mathfrak{W}}_{j}}(W_{j})$ for $i\neq j$, then by the observation above $W\in(W_{i}+{\mathfrak{W}}_{i})\cap(W_{j}+{\mathfrak{W}}_{j})=W_{i}+W_{j}+{\mathfrak{W}}_{ij}$. Therefore, $W$ divides also $P_{{\mathfrak{W}}_{ij}}(W_{i}+W_{j})$. Combining the additivity of $P_{{\mathfrak{W}}_{ij}}$ with Lemma A.5, we have
@@
which implies that $W$ divides $V_{i}+V_{j}$. As a result, every solution $\gamma$ to (4.27) satisfies
@@
This ensures that $\gamma\equiv V_{i}\;\mod P_{{\mathfrak{W}}_{i}}(W_{i})$ for every $i\in\{1,\dots,n\}$.
@@
In summary, for every choice of the polynomials $W_{1},\dots,W_{n}$, and of a solution of (4.27), this construction gives rise to a unique $n$-uple of polynomials $U_{1},\dots,U_{n}$ satisfying Proposition 4.25. Since the nonzero elements of ${\mathfrak{W}}$ are all of the same degree, and $R$ is fixed, then also the resulting $Q_{i}=U_{i}^{2}+XV_{i}^{2}$ are such that all the nonzero polynomials in $\langle Q_{1},\dots,Q_{n}\rangle_{{\mathbb{F}}_{2}}$ are all of the same degree, denoted by $\lambda$. The $n$-uple $\underline{Q}$ satisfies the conditions of Proposition 4.23 and then gives rise to a basis of a space $L_{\lambda 2^{n-1},n}$. ∎
@@
where, for every $i=1,\dots,n$, $w_{i}$ is the leading term of $W_{i}$. Then the construction of Theorem 4.26 associates with every element of $(W_{1},\dots,W_{n},R)\in{\mathcal{W}}_{d}\times k[X]$ a $n-uple$ of polynomials $(Q_{1},\dots,Q_{n})$ giving rise to a basis of a space $L_{\lambda 2^{n-1},n}$. More precisely, the construction yields $Q_{i}=U_{i}^{2}+XV_{i}^{2}$ with $V_{i}=\Delta_{n-1}(\hat{\underline{W_{i}}})$ and $U_{i}=(\alpha+R)V_{i}+\frac{V_{i}^{2}}{\Delta_{n}(\underline{W})}$, where $\alpha$ denotes the unique proper rational function (i.e. of the form $\frac{\gamma}{\Delta_{n}(\underline{W})}$ with $\gamma\in k[X]$ such that $\deg(\gamma)<\deg(\Delta_{n}(\underline{W}))$) satisfying conditions $(i)$ and $(ii)$ in Theorem 4.26.
@@
Let $\underline{Q}=(Q_{1},\dots,Q_{n})$ be the $n$-uple arising from $(W_{1},\dots,W_{n},R)$ and $\underline{Q}^{\prime}=(Q_{1}^{\prime},\dots,Q_{n}^{\prime})$ be the $n$-uple arising from $(W_{1}^{\prime},\dots,W_{n}^{\prime},R^{\prime})$. We recall that the writings
@@
are unique, and call $\underline{U},\underline{U}^{\prime},\underline{V},\underline{V}^{\prime}$ the $n$-uples arising from these. We now prove separately the statements of the corollary:

### v2

@@ 4.2.1. Many new examples of spaces $L_{\lambda 2^{n-1},n}$
We need to show that, if $\underline{U}$ and $\underline{V}$ satisfy (4.23), then $\underline{Q}$ satisfy the condition of Proposition 4.20. This latter is equivalent to $\det\left(\underline{V}^{2},\underline{U}^{2},\underline{Q^{2}},\dots,\underline{Q}^{2^{n-2}}\right)=1$, which is in turn equivalent to
@@
If the system (4.23) is satisfied, then the non-constant coefficients of the polynomial above vanish and the constant term is equal to 1. Hence we can apply Proposition 4.20 to get that $\Omega$ is a space $L_{\lambda 2^{n-1},n}$. ∎
@@
  • We then check the equations in the second line of (4.23): for each $(\epsilon_{3},\dots,\epsilon_{n})\in{\mathbb{F}}_{2}^{n-2}-\{0\}$ we let $k\in\{3,\dots,n\}$ be the smallest number such that $\epsilon_{k}\neq 0$. Since we need every $\epsilon_{i}$ with $i<k$ to be equal to 0, the vectors $\left(\underline{U},\underline{V},((1+\epsilon_{3})\underline{U}+\epsilon_{3}\underline{V})^{2},\dots,((1+\epsilon_{k})\underline{U}+\epsilon_{k}\underline{V})^{2^{k-2}}\right)$ forming the first $k$ columns inside the determinant can be rewritten as $\left(\underline{U},\underline{V},\underline{U}^{2},\dots,\underline{U}^{2^{k-3}},\underline{V}^{2^{k-2}}\right)$. By condition $(i)$, we have that all these column vectors belong to the k1k-1 dimensional space generated by $\underline{V},\dots,\underline{V}^{2^{k-2}}$ and hence
    $\det\left(\underline{U},\underline{V},\underline{U}^{2},\dots,\underline{U}^{2^{k-3}},\underline{V}^{2^{k-2}},\dots,((1+\epsilon_{n})\underline{U}+\epsilon_{n}\underline{V})^{2^{n-2}}\right)=0.$
@@
[…]
@@
    The $n$-tuple $\underline{Q}=(Q_{1},\dots,Q_{n})\in k[X]^{n}$ gives rise to a basis of a space $L_{2^{n-1}\lambda,n}$.
@@
[…]
@@
We claim that this system is equivalent to the condition that $P_{{\mathfrak{W}}_{i}}(W_{i})$ divides $\gamma+V_{i}$ for every $i\in\{1,\dots,n\}$. We start by proving the following statement:
[…]
To prove ()(*), let $i\neq j$ and set ${\mathfrak{W}}_{ij}:={\mathfrak{W}}_{i}\cap{\mathfrak{W}}_{j}$. If $W\in\langle W_{1},\dots,W_{n}\rangle-\{0\}$ is a common divisor of $P_{{\mathfrak{W}}_{i}}(W_{i})$ and $P_{{\mathfrak{W}}_{j}}(W_{j})$ for $i\neq j$, then by the observation above $W\in(W_{i}+{\mathfrak{W}}_{i})\cap(W_{j}+{\mathfrak{W}}_{j})=W_{i}+W_{j}+{\mathfrak{W}}_{ij}$. Therefore, $W$ divides also $P_{{\mathfrak{W}}_{ij}}(W_{i}+W_{j})$. Combining the additivity of $P_{{\mathfrak{W}}_{ij}}$ with Lemma A.5, we have
@@
which implies that $W$ divides $V_{i}+V_{j}$.
We choose now $\gamma$ verifying (4.26) and $W$ dividing $P_{{\mathfrak{W}}_{j}}(W_{j})$. We have $W=W_{j}+Z_{j}$ with $Z_{j}\in{\mathfrak{W}}_{j}$ and therefore $k_{W}\geq j$. If $k_{W}=j$ then we have that $\gamma\equiv V_{j}\;\mod W$ by (4.26). If $k_{W}>j$ then set $i=k_{W}$. We have that $W=W_{i}+Z_{i}$ with $Z\in{\mathfrak{W}}_{i}$ and therefore $W$ divides $P_{{\mathfrak{W}}_{i}}(W_{i})$. Using ()(*), we have that $W$ divides $V_{i}+V_{j}$ and by (4.26) we conclude that $\gamma\cong V_{i}\mod W$ and therefore $\gamma\cong V_{j}\mod W$. This shows that every solution $\gamma$ to (4.26) satisfies
@@
As a result, $\gamma\equiv V_{i}\;\mod P_{{\mathfrak{W}}_{i}}(W_{i})$ for every $i\in\{1,\dots,n\}$.
@@
In summary, for every choice of the polynomials $W_{1},\dots,W_{n}$, and of a solution of (4.26), this construction gives rise to a unique $n$-tuple of polynomials $U_{1},\dots,U_{n}$ satisfying Proposition 4.24. Since the nonzero elements of $\mathfrak{W}$} are all of the same degree, and $R$ is fixed, then also the resulting $Q_{i}=U_{i}^{2}+XV_{i}^{2}$ are such that all the nonzero polynomials in $\langle Q_{1},\dots,Q_{n}\rangle_{{\mathbb{F}}_{2}}$ are all of the same degree, denoted by $\lambda$. The $n$-tuple $\underline{Q}$ satisfies the conditions of Proposition 4.22 and then gives rise to a basis of a space $L_{\lambda 2^{n-1},n}$. ∎
@@
[…]
@@
    $\langle W_{1}^{\prime}(X),\dots,W_{n}^{\prime}(X)\rangle_{{\mathbb{F}}_{2}}=\langle W_{1}(X+b^{2}),\dots,W_{n}(X+b^{2})\rangle_{{\mathbb{F}}_{2}}\;\mbox{and}\;R^{\prime}(X)=R(X)+b.$
@@
Let $\underline{Q}=(Q_{1},\dots,Q_{n})$ be the $n$-tuple arising from $(W_{1},\dots,W_{n},R)$ and $\underline{Q}^{\prime}=(Q_{1}^{\prime},\dots,Q_{n}^{\prime})$ be the $n$-tuple arising from $(W_{1}^{\prime},\dots,W_{n}^{\prime},R^{\prime})$. We recall that the writings
@@
are unique, and call $\underline{U},\underline{U}^{\prime},\underline{V},\underline{V}^{\prime}$ the $n$-tuples arising from these. We now prove separately the statements of the corollary: